# GeneMAN: Generalizable Single-Image 3D Human Reconstruction from Multi-Source Human Data

### v2

@@ 4.2 GeneMAN Prior Models
4.3 Geometry Initialization &\&& Sculpting
As illustrated in Fig. 3, we adopt hybrid representations, incorporating NeRF [36] and DMTet [52] to reconstruct detailed human geometry from a reference image. Instead of relying on SMPL [30] for initialization, we first train the NeRF network using GeneMAN 2D and 3D prior models to craft a template-free initialized geometry. We leverage Instant-NGP [37] as our NeRF implementation due to its fast inference and ability to recover complex geometry. To supervise the reference view reconstruction, we maximize the similarity between the rendered image and the reference image $\bm{\hat{I}}$ using the following loss:
$\mathcal{L}_{\textnormal{ref}}=||\bm{\hat{m}}\odot(\bm{\hat{I}}-g(\theta;\hat{c}))||_{2}+||\bm{\hat{m}}-g_{m}(\theta;\hat{c})||_{2},$ (4)
where $g(\theta;\hat{c})$ represents the rendered image at the reference viewpoint $\hat{c}$, $\theta$ denotes the NeRF parameters and $\odot$ indicates the Hadamard product. The foreground mask is denoted as $\bm{\hat{m}}$, and $g_{m}(\theta;\hat{c})$ renders the silhouette. In addition to supervising the RGB image and mask, we utilize the geometry prior inferred from the reference image, specifically depth and normal information. The reference depth $\bm{\hat{I}_{d}}$ and normal $\bm{\hat{I}_{n}}$ are derived using the human foundation model Sapiens [21]. To enforce consistency between the rendered and reference values, we impose depth and normal losses, $\mathcal{L}_{\textnormal{depth}}$ and $\mathcal{L}_{\textnormal{normal}}$. The normal loss is calculated as the mean squared error (MSE) between the reference normal $\bm{\hat{I}_{n}}$ and $\bm{I_{n}}$, i.e., $\mathcal{L}_{\textnormal{normal}}=||\bm{\hat{I}_{n}}-\bm{I_{n}}||_{2}$. For the depth loss, we employ the normalized negative Pearson correlation:
@@ 4.3 Geometry Initialization &\&& Sculpting
To further refine the geometric details, we adopt DMTet [52], a hybrid SDF-Mesh representation that enables memory-efficient, high-resolution 3D shape reconstruction. The trained NeRF [36] is converted into a mesh, which serves as the geometric initialization for DMTet optimization. During optimization, we compute the reconstruction loss by applying both the MSE loss and the perceptual loss [54] between the rendered normal from the reference view and the reference normal $\bm{\hat{I}_{n}}$. Inspired by HumanNorm [17], we leverage its pretrained normal-adapted and depth-adapted diffusion models for novel view guidance to craft high-quality geometry. Furthermore, the SDF loss $\mathcal{L}_{\textnormal{sdf}}$ is incorporated to regularize the geometry, preventing significant deviation from the initial mesh and avoiding improper shapes. Through geometry initialization and sculpting, we achieve high-fidelity 3D human geometry with plausible poses and intricate geometric details.
Table 1: Quantitative Comparison with State-of-the-art Methods. The evaluation images are selected from in-the-wild images found on the Internet and the CAPE dataset [33]. The best results are highlighted in bold. The second-place results are underlined.
in-the-wild CAPE
@@
PIFu 26.968 0.035 0.594 26.912 0.028 0.764
GTA 25.060 0.064 0.568 30.376 0.019 0.785
TeCH 25.740 0.053 0.713 27.601 0.025 0.826
SiTH 20.412 0.129 0.608 21.992 0.048 0.815
GeneMAN (Ours) 32.238 0.013 0.730 28.490 0.015 0.838
4.4 Multi-Space Texture Refinement
With the detailed geometry in place, we implement multi-space texture refinement in both latent and pixel spaces, ensuring that the textures remain consistent and realistic. First, we perform latent space optimization using the hybrid SDS loss [41, 28], leveraging our GeneMAN 2D and 3D prior models to create an initial coarse texture. The GeneMAN 2D prior model guarantees texture realism, while the GeneMAN 3D prior model excels at maintaining view-consistent textures. We also supervise the reference view by applying an MSE loss between the reference image $\bm{\hat{I}}$ and the rendered image $g_{c}(\theta;\hat{c}))$ at viewpoint $\hat{c}$. The overall loss function for latent space optimization is formulated as:
$\mathcal{L}_{\textnormal{coarse}}$ $=\lambda^{c}_{ref}(||\bm{\hat{I}}-g_{c}(\theta;\hat{c}))||_{2}+||\bm{\hat{m}}-g_{c}(\theta;\hat{c}))||_{2})$
$+\lambda^{c}_{guid}\mathcal{L}^{c}_{\textnormal{guid}},$ (7)
Multi-View Training. Multi-view training strategies effectively enhance consistency across views [28, 53], but they typically require retraining diffusion models. In contrast, we employ a training-free approach by adding identical Gaussian noise to a batch of images rendered from different views. These images are then concatenated into a single image for inference, ensuring that the resulting multi-view images exhibit consistent textures and colors.
The coarse texture stage prioritizes learning consistent and plausible texture but still lacks realism. To enhance texture realism, we employ pixel-space optimization by minimizing the distance between the rendered multi-view images and their refined counterparts, thereby improving the UV texture map. Akin to DreamGaussian [59], we adopt the image-to-image synthesis framework of SDEdit [35] to generate the refined images. Specifically, we render the coarse image $\bm{I_{\textnormal{corase}}}$ from an arbitrary camera view $c$, perturb it with random Gaussian noise, and apply a multi-step denoising process using the GeneMAN 2D prior based ControlNet [68] $\phi_{2d}(\cdot)$ to obtain the refined image $\bm{I_{\textnormal{fine}}}$:
where $\epsilon(t_{\textnormal{start}})$ represents random noise at timestep $t_{\textnormal{start}}$, and $e$ denotes the conditioned text embeddings. We employ the MSE loss for pixel-wise reconstruction, supplemented by the LPIPS [69] loss to enhance texture details and style similarity. The total loss for pixel space optimization $\mathcal{L}_{\textnormal{fine}}$ is:
$\mathcal{L}_{\textnormal{fine}}=||\bm{I_{\textnormal{fine}}}-\bm{I_{\textnormal{coarse}}}||_{2}+\lambda_{LP}\textnormal{LPIPS}(\bm{I_{\textnormal{fine}}},\bm{I_{\textnormal{coarse}}}),$ (9)
where $\lambda_{LP}$ represents the weight of the LPIPS loss.
@@ 5.1 Implementation Details
Training Details. Our framework is built upon the open-source project ThreeStudio [10]. During the geometry stage, we progressively increase the resolution of NeRF [36] from $256$ to $384$ over $5,000$ steps. We then convert it to an explicit mesh, which serves as the geometry initialization for DMTet [52] at a resolution of $512$. We subsequently optimize DMTet for $3,000$ steps to sculpt fine-grained geometric details. In the texture stage, we perform an initial coarse texture optimization over $10,000$ steps, followed by a refinement of the texture UV map for $1,000$ steps. The full optimization process takes approximately $1.4$ hours. More details are placed in the supplementary.
@@
Testing Details. For both qualitative and quantitative evaluation, we randomly select $50$ samples from the Internet and CAPE [33]. The samples sourced from the Internet include challenging scenarios, such as humans with varying body proportions (head shots, half-body shots, full-body shots), as well as humans in diverse poses, clothing, and with personal belongings. We compare our model with state-of-the-art image-to-3D human reconstruction methods, including PIFu [47], GTA [70], TeCH [18] and SiTH [13]. For each method, we render 120 viewpoints across 360 degrees of the reconstructed results for evaluation. We compute PSNR and LPIPS [69] between the reference view and the rendered front view to evaluate reconstruction quality. Additionally, CLIP similarity [43] between the reference view and $119$ novel views is measured to assess multi-view consistency.
Figure 5: Qualitative GeneMAN Results with Complex Poses.
@@
Figure 7: Qualitative Comparison with State-of-the-art Methods on CAPE [33]. Without accurate HPS results, template-based methods [71, 13, 18] suffer from artifacts like the “bent-leg” effect. In contrast, GeneMAN successfully reconstructs 3D humans with natural poses, detailed geometry, and realistic textures.
5.2 Quantitative Comparison
Tab. 1 provides quantitative metrics of GeneMAN with baseline methods on in-the-wild images and CAPE [33]. For in-the-wild images, GeneMAN outperforms all competing approaches, showcasing its superiority in generalizable human reconstruction, with the highest CLIP-Similarity [43] indicating better multi-view consistency. It also leads in PSNR and LPIPS, highlighting its improved reconstruction quality. On CAPE, GeneMAN continues to excel, achieving the best results in LPIPS and CLIP-Similarity while maintaining competitive performance in PSNR, proving its effectiveness on laboratory dataset as well.
Figure 8: Ablation on Geometry Reconstruction and Texture Generation. (a) shows the geometry initialization, (b) recovers finer details through geometry sculpting, (c) applies latent-space texturing, and (d) achieves the final reconstruction with pixel-space optimization.
Figure 9: Ablation on GeneMAN 2D and 3D Prior Models. (a) and (b) demonstrate the 2D prior model improves multi-view consistency and texture details, while (c) and (d) illustrate enhanced pose naturalness and geometric accuracy achieved by the 3D prior model.
Figure 10: User Study. Across all examples, $73.09\%$ of subjects prefer our method over the baselines in terms of geometry and texture, highlighting the significant superiority of our approach in recovering accurate geometry and high-fidelity appearances.
5.3 Qualitative Comparison
We present qualitative results on testing samples from in-the-wild images and the CAPE [33] dataset, showcased in Fig. 6 and Fig. 7, respectively. To better illustrate the reconstruction quality, we visualize the multi-view surface normals and color renderings. PIFu [47] fails to recover realistic geometry and appearance, particularly in side views. Template-based approaches, including GTA [71], TeCH [18] and SiTH [13] heavily rely on precise human pose and shape estimation (HPS). However, HPS methods [9] often produce artifacts such as “bent legs” and struggle with the shapes of children, as seen in both Fig. 6 and Fig. 7. These limitations lead to cumulative errors in geometry reconstruction that are difficult to eliminate, even with refinement. While TeCH captures more lifelike details with optimization-based refinement, the reconstructed surface appears excessively noisy, and the textures are inconsistent, especially in the back views of garments. In contrast, GeneMAN demonstrates exceptional generalizability to in-the-wild images featuring diverse clothing styles, poses and varying body proportions. Its template-free design enables superior modeling of personal belongings and loose clothing, such as basketball, skirts and dresses, while maintaining detailed, realistic geometry and natural body shapes. GeneMAN also produces high-fidelity, multi-view consistent textures for clothing and hairstyles. In Fig. 5, we present additional results to highlight the robust reconstruction capabilities of GeneMAN under complex poses. More results on complex poses are provided in the supplementary.
@@ 5.5 Ablation Study
Effectiveness of Geometry Reconstruction and Texture Generation. Fig. 10 analyzes the effectiveness of each stage proposed in our framework. Compared with geometry initialization (a), geometry sculpting in (b) smooths the excessively noisy surface and recovers high-frequency details, such as clothing wrinkles and facial features. In (c), the latent space texturing yields a reasonably satisfactory result; however, the back view remains inconsistent, and the texture appears slightly blurred. To address this, (d) enhance the coarse texture into a final refined texture through pixel space texturing, optimizing the UV map.
Effectiveness of GeneMAN 2D Prior Model. To evaluate the efficacy of our GeneMAN 2D prior model, we compare the mesh results extracted from the NeRF [36] stage utilizing the pretrained DeepFloyd-IF [46] model (“original 2D prior”) versus our GeneMAN 2D prior model (“GeneMAN 2D prior”). Note that we select the NeRF stage output for visualization to clearly demonstrate the impact of the diffusion prior, whereas subsequent geometry and texture disentanglement may obscure this effect. As illustrated in Fig. 10 (a) and (b), the red-boxed regions reveal that the original 2D prior results in an uneven shirt hem, with the back extending lower than the front. In contrast, GeneMAN 2D prior model maintains multi-view consistency in both geometry and texture, enhancing overall reconstruction.
Effectiveness of GeneMAN 3D Prior Model. To assess how GeneMAN 3D prior model (“GeneMAN 3D prior”) make influence on reconstruction results, we conduct a comparison by substituting it with the original Zero-1-to-3 [28] model (“original 3D prior”) in our framework. As indicated in Fig. 10, (c) reveals unnatural human poses, with the neck and head leaning forward, misaligned with the input image. In contrast, our GeneMAN 3D prior model, finetuned on a large-scale multi-source human dataset, captures a more natural pose and shape prior, as shown in (d). This comparison highlights that our GeneMAN 3D prior offers more reliable and coherent guidance, leading to improved geometric accuracy and a more realistic human structure.
@@ 6 Conclusion
[…]
Supplementary Material

### v3

@@ 4.2 GeneMAN Prior Models
Figure 3: Geometry Initialization &\& Sculpting. During the geometry reconstruction stage, we initialize a template-free geometry using NeRF Mildenhall et al. (2021), incorporating GeneMAN 2D and 3D priors with SDS losses. Alongside diffusion-based guidance, a reference loss ensures alignment with the input image. We then convert NeRF into DMTet Shen et al. (2021) for high-resolution refinement, guided by pretrained human-specific normal- and depth-adapted diffusion models Huang et al. (2024a).
4.3 Geometry Initialization &\& Sculpting
As illustrated in Fig. 3, we adopt hybrid representations, incorporating NeRF Mildenhall et al. (2021) and DMTet Shen et al. (2021) to reconstruct detailed human geometry from a reference image. Instead of relying on SMPL Loper et al. (2023) for initialization, we first train the NeRF network Mildenhall et al. (2021) using GeneMAN 2D and 3D prior models to craft a template-free initialized geometry. We leverage Instant-NGP Müller et al. (2022) as our NeRF implementation due to its fast inference and ability to recover complex geometry. To supervise the reference view reconstruction, we maximize the similarity between the rendered image and the reference image $\bm{\hat{I}}$ using the following loss:
[…]
@@ 4.3 Geometry Initialization &\& Sculpting
To further refine the geometric details, we adopt DMTet Shen et al. (2021), a hybrid SDF-Mesh representation that enables memory-efficient, high-resolution 3D shape reconstruction. The trained NeRF Mildenhall et al. (2021) is converted into a mesh, which serves as the geometric initialization for DMTet optimization. During optimization, we compute the reconstruction loss by applying both the MSE loss and the perceptual loss Simonyan (2014) between the rendered normal from the reference view and the reference normal $\bm{\hat{I}_{n}}$. Inspired by HumanNorm Huang et al. (2024a), we leverage its pretrained normal-adapted and depth-adapted diffusion models for novel view guidance to craft high-quality geometry. Furthermore, the SDF loss $\mathcal{L}_{\textnormal{sdf}}$ is incorporated to regularize the geometry, preventing significant deviation from the initial mesh and avoiding improper shapes. Through geometry initialization and sculpting, we achieve high-fidelity 3D human geometry with plausible poses and intricate geometric details.
4.4 Multi-Space Texture Refinement
With the detailed geometry in place, we implement multi-space texture refinement in both latent and pixel spaces, ensuring that the textures remain consistent and realistic. First, we perform latent space optimization using the hybrid SDS loss Poole et al. (2022); Liu et al. (2023), leveraging our GeneMAN 2D and 3D prior models to create an initial coarse texture. The GeneMAN 2D prior model guarantees texture realism, while the GeneMAN 3D prior model excels at maintaining view-consistent textures. We also supervise the reference view by applying an MSE loss between the reference image $\bm{\hat{I}}$ and the rendered image $)g_{c}(\theta;\hat{c}$)) at viewpoint $\hat{c}$. The overall loss function for latent space optimization is formulated as:
$\mathcal{L}_{\textnormal{coarse}}$ $=\lambda^{c}_{ref}(||\bm{\hat{I}}-g_{c}(\theta;\hat{c}))||_{2}+||\bm{\hat{m}}-g_{c}(\theta;\hat{c}))||_{2})+\lambda^{c}_{guid}\mathcal{L}^{c}_{\textnormal{guid}},$ (7)
Multi-View Training. Multi-view training strategies effectively enhance consistency across views Liu et al. (2023); Shi et al. (2023), but they typically require retraining diffusion models. In contrast, we employ a training-free approach by adding identical Gaussian noise to a batch of images rendered from different views. These images are then concatenated into a single image for inference, ensuring that the resulting multi-view images exhibit consistent textures and colors.
The coarse texture stage prioritizes learning consistent and plausible texture but still lacks realism. To enhance texture realism, we employ pixel-space optimization by minimizing the distance between the rendered multi-view images and their refined counterparts, thereby improving the UV texture map. Akin to DreamGaussian Tang et al. (2023), we adopt the image-to-image synthesis framework of SDEdit Meng et al. (2021) to generate the refined images. Specifically, we render the coarse image $\bm{I_{\textnormal{corase}}}$ from an arbitrary camera view $c$, perturb it with random Gaussian noise, and apply a multi-step denoising process using the GeneMAN 2D prior based ControlNet Zhang et al. (2023) $\phi_{2d}(\cdot)$ to obtain the refined image $\bm{I_{\textnormal{fine}}}$:
where $\epsilon(t_{\textnormal{start}})$ represents random noise at timestep $t_{\textnormal{start}}$, and $e$ denotes the conditioned text embeddings. We employ MSE loss for pixel-wise reconstruction and LPIPS Zhang et al. (2018) loss to enhance texture details. Let $\lambda_{LP}$ denote the weight of the LPIPS loss. The total loss for pixel space optimization $\mathcal{L}_{\textnormal{fine}}$ is:
@@ 5.1 Implementation Details
Training Details. Our framework is built upon the open-source project ThreeStudio Guo et al. (2023). During the geometry stage, we progressively increase the resolution of NeRF Mildenhall et al. (2021) from $256$ to $384$ over $5,000$ steps. We then convert it to an explicit mesh, which serves as the geometry initialization for DMTet Shen et al. (2021) at a resolution of $512$. We subsequently optimize DMTet for $3,000$ steps to sculpt fine-grained geometric details. In the texture stage, we perform an initial coarse texture optimization over $10,000$ steps, followed by a refinement of the texture UV map for $1,000$ steps. The full optimization process takes approximately $1.4$ hours on single NVIDIA A100 80G GPU. Additional details are provided in the supplementary materials.
@@
Testing Details. For both qualitative and quantitative evaluation, we randomly select $50$ samples from the Internet and CAPE Ma et al. (2020). The samples sourced from the Internet include challenging scenarios, such as humans with varying body proportions (head shots, half-body shots, full-body shots), as well as humans in diverse poses, clothing, and with personal belongings. We compare our model with state-of-the-art image-to-3D human reconstruction methods, including PIFu Saito et al. (2019), GTA Zhang et al. (2024a), TeCH Huang et al. (2024b) and SiTH Ho et al. (2024). For each method, we render 120 viewpoints across 360 degrees of the reconstructed results for evaluation. We compute PSNR and LPIPS Zhang et al. (2018) between the reference view and the rendered front view to evaluate reconstruction quality. Additionally, CLIP similarity Radford et al. (2021) between the reference view and $119$ novel views is measured to assess multi-view consistency.
Table 1: Quantitative Comparison with State-of-the-art Methods. Evaluation images are sourced from in-the-wild and the CAPE dataset Ma et al. (2020). Best results are in bold, second-best are underlined.
Method in-the-wild CAPE
[…]
PIFu Saito et al. (2019) 26.968 0.035 0.594 26.912 0.028 0.764
GTA Zhang et al. (2024b) 25.060 0.064 0.568 30.376 0.019 0.785
TeCH Huang et al. (2024b) 25.740 0.053 0.713 27.601 0.025 0.826
SiTH Ho et al. (2024) 20.412 0.129 0.608 21.992 0.048 0.815
GeneMAN (Ours) 32.238 0.013 0.730 28.490 0.015 0.838
5.2 Quantitative Comparison
Tab. 1 presents quantitative metrics comparing GeneMAN with baseline methods on in-the-wild images and the CAPE dataset Ma et al. (2020). For in-the-wild images, GeneMAN outperforms all competing approaches, showcasing its superiority in generalizable human reconstruction, with the highest CLIP-Similarity Radford et al. (2021) indicating better multi-view consistency. It also leads in PSNR and LPIPS, highlighting its improved reconstruction quality. On CAPE dataset, GeneMAN continues to excel, achieving the best results in LPIPS and CLIP-Similarity while maintaining competitive performance in PSNR, proving its effectiveness on laboratory dataset as well. Additional geometric results are placed in the supplementary materials.
5.3 Qualitative Comparison
Figure 5: Qualitative Results of GeneMANwith Complex Poses.
We present qualitative results on testing samples from in-the-wild images and the CAPE Ma et al. (2020) dataset, showcased in Fig. 6 and Fig. 7, respectively. To better illustrate the reconstruction quality, we visualize the multi-view surface normals and color renderings. PIFu Saito et al. (2019) fails to recover realistic geometry and appearance, particularly in side views. Template-based approaches, including GTA Zhang et al. (2024b), TeCH Huang et al. (2024b) and SiTH Ho et al. (2024) heavily rely on precise human pose and shape estimation (HPS). However, HPS methods Feng et al. (2021) often produce artifacts such as “bent legs” and struggle with the shapes of children, as seen in both Fig. 6 and Fig. 7. These limitations lead to cumulative errors in geometry reconstruction that are difficult to eliminate, even with refinement. While TeCH captures more lifelike details with optimization-based refinement, the reconstructed surface appears excessively noisy, and the textures are inconsistent, especially in the back views of garments. In contrast, GeneMAN demonstrates exceptional generalizability to in-the-wild images featuring diverse clothing styles, poses and varying body proportions. Its template-free design enables superior modeling of personal belongings and loose clothing, such as basketball, skirts and dresses, while maintaining detailed, realistic geometry and natural body shapes. GeneMAN also produces high-fidelity, multi-view consistent textures for clothing and hairstyles. In Fig. 5, we present additional results to highlight the robust reconstruction capabilities of GeneMAN under complex poses. More results on complex poses are provided in the supplementary.
@@ 5.3 Qualitative Comparison
Figure 7: Qualitative Comparison with State-of-the-art Methods on CAPE Ma et al. (2020). Without accurate HPS results, template-based methods Zhang et al. (2024b); Ho et al. (2024); Huang et al. (2024b) suffer from artifacts like the “bent-leg” effect. In contrast, GeneMAN reconstructs humans with natural poses, detailed geometry, and realistic textures.
@@ 5.4 User Study
Figure 8: Ablation on Geometry and Texture Reconstruction. (a) shows the geometry initialization, (b) recovers finer details through geometry sculpting, (c) applies latent-space texturing, and (d) achieves the final reconstruction with pixel-space optimization.
Figure 9: Ablation on GeneMAN 2D and 3D Prior Models. (a) and (b) demonstrate the 2D prior model improves multi-view consistency and texture details. (c) and (d) illustrate the enhanced pose naturalness and geometric accuracy from the 3D prior model.
Figure 10: User Study. $73.09\%$ of subjects prefer our method over the baselines in terms of geometry and texture, highlighting the significant superiority of our approach in recovering accurate geometry and high-fidelity appearances.
@@ 5.5 Ablation Study
Effectiveness of Geometry and Texture Reconstruction. Fig. 10 analyzes the effectiveness of each stage proposed in our framework. Compared with geometry initialization (a), geometry sculpting in (b) smooths the excessively noisy surface and recovers high-frequency details, such as clothing wrinkles and facial features. In (c), the latent space texturing yields a reasonably satisfactory result; however, the back view remains inconsistent, and the texture appears slightly blurred. To address this, (d) enhance the coarse texture into a final refined texture through pixel space texturing.
Effectiveness of GeneMAN 2D Prior Model. To assess the efficacy of GeneMAN 2D prior model, we compare the mesh results extracted from the NeRF Mildenhall et al. (2021) stage utilizing the pretrained DeepFloyd-IF Saharia et al. (2022b) model (“original 2D prior”) versus GeneMAN 2D prior model (“GeneMAN 2D prior”). Note that we select the NeRF stage output for visualization to clearly demonstrate the impact of the diffusion prior, whereas subsequent geometry and texture disentanglement may obscure this effect. As shown in Fig. 10 (a) and (b), the red-boxed regions reveal that the original 2D prior results in an uneven shirt hem, with the back extending lower than the front. In contrast, GeneMAN 2D prior model ensures multi-view consistency in both geometry and texture, enhancing overall reconstruction.
Effectiveness of GeneMAN 3D Prior Model. To assess how GeneMAN 3D prior model (“GeneMAN 3D prior”) make influence on reconstruction results, we conduct a comparison by substituting it with the original Zero-1-to-3 Liu et al. (2023) model (“original 3D prior”) in our framework. As indicated in Fig. 10, (c) reveals unnatural human poses, with the neck and head leaning forward, misaligned with the input image. In contrast, our GeneMAN 3D prior model, finetuned on a large-scale multi-source human dataset, captures a more natural pose and shape prior, as shown in (d). This comparison highlights that our GeneMAN 3D prior offers more reliable and coherent guidance, leading to improved geometric accuracy and a more realistic human structure.
@@ 6 Conclusion
Acknowledgments
This work was partly supported by National Natural Science Foundation of China (62506229, 62302295), and Natural Science Foundation of Shanghai (25ZR1402268, 2021SHZDZX0102). We sincerely thank Wayne Wu, Shikai Li, and Honglin He for their insightful suggestions and feedback on the manuscript.
Supplementary Materials